# DAWN-SI: Data-aware and Noise-Informed Stochastic Interpolation for Solving Inverse Problems

### v1

@@ 3 Numerical Experiments
In this section, we experiment with our method on a few common datasets and two broadly applicable inverse problems: image deblurring and tomography (see Appendix A for details). We provide additional information on the experimental settings, hyperparameter choices, and network architectures in Tables 4, 5 and 3 for image deblurring task and Table 6 for tomography task in Appendix C.
Training methodology. For our experiments, we considered two types of data-aware velocity estimator networks: (i) DAW-SI: the estimator with no noise-embedding, and (ii) DAWN-SI: the estimator with trainable noise-embedding. We employed antithetic sampling during training. Starting with an input batch of clean images ${\bf x}_{1}^{\prime}$, an ${\bf x}_{0}^{\prime}\sim\mathcal{N}(0,{\bf I})$ was sampled and antithetic pairs ${\bf x}_{1}$ and ${\bf x}_{0}$,
@@
were generated by concatenation along the batch dimension. For a large number of samples, the sample mean of independent random variables converges to the true mean. However, the convergence can be slow due to the variance of the estimator being large. Antithetic sampling helped reduce this variance by generating pairs of negatively correlated samples, thereby improving convergence. The data was generated by employing the forward model along with a Gaussian noise injection, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$, where ${\bf z}\sim\mathcal{N}(0,{\bf I})$. The value of $\sigma$ was set to $p\%$ of the range of values in the data ${\bf b}$, where $p$ was sampled uniformly in (0, 20). ${\bf x}_{t}$ and ${\bf v}$ were computed from ${\bf x}_{1}$ and ${\bf x}_{0}$ following Equation 1 and Equation 2. Using the transformation $E_{\boldsymbol{\eta}}(f({\bf b}))=E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}$), the predicted velocity of the estimator was $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-SI estimator and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-SI estimator. Here, $E_{\boldsymbol{\eta}}$ was chosen to be a single-layer convolutional neural network. The loss for an epoch was computed as given in Equation 14.
@@
Inference. For inference on a noisy data, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$ for some image ${\bf x}_{1}$, we start with a randomly sampled ${\bf x}_{0}\sim\mathcal{N}(0,{\bf I})$ at $t=0$ and perform Fourth-Order Runga Kutta numerical integration to solve the ODE in Equation 10 with a step size $h=1/100$, where the velocity was computed using the trained estimator as $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-SI estimator and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-SI estimator at time $t$. For each image ${\bf x}_{1}$, the inference was run 32 times starting from 32 different realizations of ${\bf x}_{0}$ and the resulting reconstructed images (which are samples from the posterior $\pi({\bf x}_{1}|{\bf b})$) were averaged to generate the final recovered image. The uncertainty at the pixel-level could also be estimated from these samples by computing their standard deviation, as shown in Figure 4.
@@
Baselines. For the image deblurring task, we compare our methods DAW-SI and DAWN-SI to diffusion-based image deblurring [Chung et al., 2022a] and InverseUNetODE [Eliasof et al., 2024]. The details for the training setup and hyperparameter choices for the baselines have been provided in Appendix C. All baselines were trained with Gaussian noise injection in the data in the same way as used for training of DAW-SI and DAWN-SI methods for fair comparison.
@@
Metrics. For performance evaluation, we compute mean squared error (MSE), misfit, structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) between the ground truth and the reconstructed image. Additional details on these metrics are provided in Appendix B.
@@ 3.1 Image Deblurring
We present the results of image deblurring task in Table 1 at noise level $p=5\%$ in blurred data for MNIST, STL10 and CIFAR10 datasets. For all metrics, DAWN-SI beats all other methods for all datasets by large margins, except for the CIFAR10 dataset, where DAW-SI performed marginally better than DAWN-SI. Moreover, on the CIFAR10 dataset, the Diffusion model fits the data the best achieving the lowest value of the misfit metric. We also performed ablation studies for different noise levels in data for both DAW-SI and DAWN-SI models, as shown in Figure 5. For all metrics, DAWN-SI was more robust to noise in data, especially for higher noise levels over all datasets with the crossing point happening at values of $p\lesssim 5\%$ for all metrics. Some example images for the deblurring task for different levels of noise in data are presented in Figures 7, 8 and 9 in Appendix D. Based on empirical evidence from training, we believe that the results for both DAW-SI and DAWN-SI could be improved further with more training epochs.
Figure 5: Assessing the sensitivity of the metrics for the recovered deblurred image as a function of the level of noise ($p\%$) added to the data. The DAWN-SI  is more robust to noise across all three datasets, MNIST, STL10 and CIFAR10 than DAW-SI, with the crossing point happening at values of $p$ $\lesssim$ $5\%$ for all metrics.
Table 1: Comparison between the performances of different methods (Diffusion, InverseUNetODE, and our methods DAW-SI and DAWN-SI) for the deblurring task for MNIST, STL10 and CIFAR10 datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for a large blurring kernel (for details, see Section A.1) at a noise level $p=5\%$ in the blurred data. Here, we report mean $\pm$ standard deviation over all images in the test set. First and second best performances are shown in bold and red, respectively.
@@
DAW-SI (Ours) 0.156 $\pm$ 0.078 0.043 $\pm$ 0.012 0.825 $\pm$ 0.062 18.899 $\pm$ 2.472
DAWN-SI (Ours) 0.073 $\pm$ 0.041 0.032 $\pm$ 0.009 0.901 $\pm$ 0.042 22.277 $\pm$ 2.612
@@
DAW-SI (Ours) 0.344 $\pm$ 0.132 0.049 $\pm$ 0.012 0.382 $\pm$ 0.064 17.434 $\pm$ 1.465
DAWN-SI (Ours) 0.113 $\pm$ 0.056 0.043 $\pm$ 0.013 0.644 $\pm$ 0.082 22.423 $\pm$ 1.904
@@
DAW-SI (Ours) 0.152 $\pm$ 0.071 0.057 $\pm$ 0.025 0.567 $\pm$ 0.077 19.765 $\pm$ 1.520
DAWN-SI (Ours) 0.156 $\pm$ 0.077 0.055 $\pm$ 0.023 0.554 $\pm$ 0.081 19.743 $\pm$ 1.712
Figure 6: An example from OrganAMNIST dataset showing the different samples (labeled ${\bf x}^{(1)}_{1}$ to ${\bf x}^{(4)}_{1}$) drawn from the posterior starting from different initial ${\bf x}_{0}$ for both DAW-SI (top) and DAWN-SI (bottom). The posterior mean $\bar{{\bf x}}_{1}$ and standard deviation ${\sigma}_{{\bar{\bf x}}_{1}}$ were computed from 32 such samples, given the sinogram data (at noise level $p=5\%$). The red box highlights a lobe, which exhibits variability in its size and shape across different samples. This variability reflects the model’s uncertainty in reconstructing (the boundary of) such anatomical features.
@@ 3.2 Tomography
We present the results of our methods DAW-SI and DAWN-SI for the tomography task in Table 2 at noise level $p=5\%$ in the sinogram data for OrganAMNIST and OrganCMNIST datasets [Yang et al., 2023]. For this task, DAWN-SI outperformed DAW-SI on all metrics for both datasets. Some example images for the tomography task for different levels of noise are presented in Figures 10 and 11 in Appendix D. In medical image analysis, accurately estimating the size and boundaries of a lobe (such as the one shown in red box in Figure 6) is crucial for diagnostics, especially in cases where its size can influence medical decisions. We quantify uncertainty for both DAW-SI and DAWN-SI by computing mean and standard deviation across 32 samples from the learned posterior. While the mean represents the most probable reconstruction, the standard deviation map highlights regions of higher variability, indicating areas of uncertainty. From Figure 6, DAWN-SI gives a robust understanding of uncertainty by highlighting boundaries of objects as regions of highest uncertainty, a feature similar to what was observed on the MNIST dataset (see Figure 4).
Table 2: Comparison between the performances of our methods (DAW-SI and DAWN-SI) for the tomography task for OrganAMNIST and OrganCMNIST datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for $N_{\text{angles}}=360$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the image dimension (for details, see Section A.2) at a noise level $p=5\%$ in the sinogram data. Here, we report mean $\pm$ standard deviation over 5000 images in the test set. Best performance is shown in bold.
@@
OrganAMNIST DAW-SI (Ours) 0.008 $\pm$ 0.002 0.041 $\pm$ 0.051 0.575 $\pm$ 0.108 20.389 $\pm$ 1.767
DAWN-SI (Ours) 0.004 $\pm$ 0.001 0.036 $\pm$ 0.060 0.713 $\pm$ 0.090 23.244 $\pm$ 1.553
OrganCMNIST DAW-SI (Ours) 0.014 $\pm$ 0.008 0.156 $\pm$ 1.424 0.511 $\pm$ 0.140 17.881 $\pm$ 2.624
DAWN-SI (Ours) 0.007 $\pm$ 0.007 0.151 $\pm$ 1.287 0.675 $\pm$ 0.116 21.630 $\pm$ 3.199
@@ 4 Conclusions
In this paper, we presented DAWN-SI, a framework for addressing highly ill-posed inverse problems by efficiently incorporating data and noise into the SI process. Our experiments showed that our proposed method consistently outperformed existing methods in tasks like image deblurring, with significant improvements in key performance metrics. The ability to sample from the learned posterior enables the exploration of the solution space and facilitates in uncertainty quantification, which is critical for real-world applications. Future work will focus on refining the model to enhance efficiency and applicability to more ill-posed problems, including integration of advanced noise modeling techniques for extreme noise conditions.
Ethics and Reproducibility Statements
Ethics Statement. In this work, we do not release any datasets or models that could be misused, and we believe our research carries no direct or indirect negative societal implications. We do not work with sensitive or privacy-related data, nor do we develop methods that could be applied to harmful purposes. To the best of our knowledge, this study raises no ethical concerns or risks of negative impact. Additionally, our research does not involve human subjects or crowdsourcing. We also confirm that there are no conflicts of interest or external sponsorships influencing the objectivity or results of this study.
Reproducibility Statement. In Section 3, we outline the training methodology employed in our experiments, while in Appendix C, we provide comprehensive supplementary information, including references to the baselines, detailed dataset descriptions, the experimental settings for each task, and the hyperparameter used in our study. All experiments presented in Section 3 were conducted on publicly available benchmarks, while the experiment in Example 2.1 was conducted on simulated data. To further facilitate the reproducibility of our work, we will release all the data and code to reproduce our empirical evaluation upon acceptance.
@@ Appendix C Experimental Settings
For the image deblurring task, the experiments were conducted on MNIST, STL10 and CIFAR10 datasets. The networks used were Diffusion model, InverseUNetODE as baselines and our proposed methods DAW-SI and DAWN-SI. The key details of the experimental setup for each of these methods are summarized in Tables 3, 4 and 5. For the tomography task, the experiments were conducted using DAW-SI and DAWN-SI methods on OrganAMNIST and OrganCMNIST datasets derived from the MedMNIST data library [Yang et al., 2023]. The key details of the experimental setup are summarized in Table 6. All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch [Paszke et al., 2017].
Table 3: Experimental details for the image deblurring task for our methods DAW-SI and DAWN-SI.
@@
DAW-SI: UNet with 3 levels with [$c$, 16, 32] filters with residual
@@
DAWN-SI: UNet with 3 level levels with [$c$, 16, 32] filters with
@@
DAW-SI: UNet with 5 levels with [3, 16, 32. 64, 128] filters
@@
DAWN-SI: UNet with 5 levels with [3, 16, 32, 64, 128] filters
@@
DAW-SI: max_epochs = 3000;
DAWN-SI: max_epochs = 3000
@@
Table 6: Experimental details for the tomography task for our methods DAW-SI and DAWN-SI.
@@
DAW-SI: UNet with 5 levels with [3, 16, 32. 64, 128] filters
@@
DAWN-SI: UNet with 5 levels with [3, 16, 32, 64, 128] filters
@@
DAW-SI: max_epochs = 3000;
DAWN-SI: max_epochs = 3000

### v2

@@ 3 Numerical Experiments
In this section, we experiment with our method on a few common datasets and two broadly applicable inverse problems: image deblurring and tomography (see A for details). We provide additional information on the experimental settings, hyperparameter choices, and network architectures in Tables 4, 5 and 3 for image deblurring task and Table 6 for tomography task in C.
Training methodology. For our experiments, we considered two types of data-aware velocity estimator networks: (i) DAW-FM: the estimator with no noise-embedding, and (ii) DAWN-FM : the estimator with trainable noise-embedding. We employed antithetic sampling during training. Starting with an input batch of clean images ${\bf x}_{1}^{\prime}$, an ${\bf x}_{0}^{\prime}\sim\mathcal{N}(0,{\bf I})$ was sampled and antithetic pairs ${\bf x}_{1}$ and ${\bf x}_{0}$,
@@
were generated by concatenation along the batch dimension. For a large number of samples, the sample mean of independent random variables converges to the true mean. However, the convergence can be slow due to the variance of the estimator being large. Antithetic sampling helped reduce this variance by generating pairs of negatively correlated samples, thereby improving convergence. The data was generated by employing the forward model along with a Gaussian noise injection, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$, where ${\bf z}\sim\mathcal{N}(0,{\bf I})$. The value of $\sigma$ was set to $p\%$ of the range of values in the data ${\bf b}$, where $p$ was sampled uniformly in (0, 20). ${\bf x}_{t}$ and ${\bf v}$ were computed from ${\bf x}_{1}$ and ${\bf x}_{0}$ following Equation 1 and Equation 2. Using the transformation $E_{\boldsymbol{\eta}}(f({\bf b}))=E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}$), the predicted velocity of the estimator was $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-FM estimator and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-FM estimator. Here, $E_{\boldsymbol{\eta}}$ was chosen to be a single-layer convolutional neural network. The loss for an epoch was computed as given in Equation 14.
@@
Inference. For inference on a noisy data, ${\bf b}={\bf A}{\bf x}_{1}+\sigma{\bf z}$ for some image ${\bf x}_{1}$, we start with a randomly sampled ${\bf x}_{0}\sim\mathcal{N}(0,{\bf I})$ at $t=0$ and perform Fourth-Order Runga Kutta numerical integration to solve the ODE in Equation 10 with a step size $h=1/100$, where the velocity was computed using the trained estimator as $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t)$ for the DAW-FM estimator and $s_{\boldsymbol{\theta}}({\bf x}_{t},E_{\boldsymbol{\eta}}({\bf A}^{\top}{\bf b}),t,\sigma)$ for the DAWN-FM estimator at time $t$. For each image ${\bf x}_{1}$, the inference was run 32 times starting from 32 different realizations of ${\bf x}_{0}$ and the resulting reconstructed images (which are samples from the posterior $\pi({\bf x}_{1}|{\bf b})$) were averaged to generate the final recovered image. The uncertainty at the pixel-level could also be estimated from these samples by computing their standard deviation, as shown in Figure 4.
@@
Baselines. For the image deblurring task, we compare our methods DAW-FM and DAWN-FM to diffusion-based image deblurring [Chung et al., 2022a] and InverseUNetODE [Eliasof et al., 2024]. The details for the training setup and hyperparameter choices for the baselines have been provided in C. All baselines were trained with Gaussian noise injection in the data in the same way as used for training of DAW-FM and DAWN-FM methods for fair comparison.
@@
Metrics. For performance evaluation, we compute mean squared error (MSE), misfit, structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) between the ground truth and the reconstructed image. Additional details on these metrics are provided in B.
@@ 3.1 Image deblurring
We present the results of image deblurring task in Table 1 at noise level $p=5\%$ in blurred data for MNIST, STL10 and CIFAR10 datasets. For all metrics, DAWN-FM beats all other methods for all datasets by large margins, except for the CIFAR10 dataset, where DAW-FM performed marginally better than DAWN-FM. Moreover, on the CIFAR10 dataset, the diffusion model fits the data the best achieving the lowest value of the misfit metric. We also performed ablation studies for different noise levels in data for both DAW-FM and DAWN-FM models, as shown in Figure 5. For all metrics, DAWN-FM was more robust to noise in data, especially for higher noise levels over all datasets with the crossing point happening at values of $p\lesssim 5\%$ for all metrics. Some example images for the deblurring task for different levels of noise in data are presented in Figures 7, 8 and 9 in D. Based on empirical evidence from training, we believe that the results for both DAW-FM and DAWN-FM could be improved further with more training epochs.
Figure 5: Assessing the sensitivity of the metrics for the recovered deblurred image as a function of the level of noise ($p\%$) added to the data. The DAWN-FM is more robust to noise across all three datasets, MNIST, STL10 and CIFAR10 than DAW-FM, with the crossing point happening at values of $p$ $\lesssim$ $5\%$ for all metrics.
Table 1: Comparison between the performances of different methods (Diffusion, InverseUNetODE, and our methods DAW-FM and DAWN-FM) for the deblurring task for MNIST, STL10 and CIFAR10 datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for a large blurring kernel (for details, see A.1) at a noise level $p=5\%$ in the blurred data. Here, we report mean $\pm$ standard deviation over all images in the test set. First and second best performances are shown in bold and red, respectively.
@@
DAW-FM (Ours) 0.156 $\pm$ 0.078 0.043 $\pm$ 0.012 0.825 $\pm$ 0.062 18.899 $\pm$ 2.472
DAWN-FM (Ours) 0.073 $\pm$ 0.041 0.032 $\pm$ 0.009 0.901 $\pm$ 0.042 22.277 $\pm$ 2.612
@@
DAW-FM (Ours) 0.344 $\pm$ 0.132 0.049 $\pm$ 0.012 0.382 $\pm$ 0.064 17.434 $\pm$ 1.465
DAWN-FM (Ours) 0.113 $\pm$ 0.056 0.043 $\pm$ 0.013 0.644 $\pm$ 0.082 22.423 $\pm$ 1.904
@@
DAW-FM (Ours) 0.152 $\pm$ 0.071 0.057 $\pm$ 0.025 0.567 $\pm$ 0.077 19.765 $\pm$ 1.520
DAWN-FM (Ours) 0.156 $\pm$ 0.077 0.055 $\pm$ 0.023 0.554 $\pm$ 0.081 19.743 $\pm$ 1.712
Figure 6: An example from OrganAMNIST dataset showing the different samples (labeled ${\bf x}^{(1)}_{1}$ to ${\bf x}^{(4)}_{1}$) drawn from the posterior starting from different initial ${\bf x}_{0}$ for both DAW-FM (top) and DAWN-FM (bottom). The posterior mean $\bar{{\bf x}}_{1}$ and standard deviation ${\sigma}_{{\bar{\bf x}}_{1}}$ were computed from 32 such samples, given the sinogram data (at noise level $p=5\%$). The red box highlights a lobe, which exhibits variability in its size and shape across different samples. This variability reflects the model’s uncertainty in reconstructing (the boundary of) such anatomical features.
@@ 3.2 Tomography
Table 2: Comparison between the performances of our methods (DAW-FM and DAWN-FM) for the tomography task for OrganAMNIST and OrganCMNIST datasets using metrics MSE, MISFIT, SSIM and PSNR. All the evaluations have been presented on the test set for $N_{\text{angles}}=360$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the image dimension (for details, see A.2) at a noise level $p=5\%$ in the sinogram data. Here, we report mean $\pm$ standard deviation over 5000 images in the test set. Best performance is shown in bold.
@@
OrganAMNIST DAW-FM (Ours) 0.008 $\pm$ 0.002 0.041 $\pm$ 0.051 0.575 $\pm$ 0.108 20.389 $\pm$ 1.767
DAWN-FM (Ours) 0.004 $\pm$ 0.001 0.036 $\pm$ 0.060 0.713 $\pm$ 0.090 23.244 $\pm$ 1.553
OrganCMNIST DAW-FM (Ours) 0.014 $\pm$ 0.008 0.156 $\pm$ 1.424 0.511 $\pm$ 0.140 17.881 $\pm$ 2.624
DAWN-FM (Ours) 0.007 $\pm$ 0.007 0.151 $\pm$ 1.287 0.675 $\pm$ 0.116 21.630 $\pm$ 3.199
We present the results of our methods DAW-FM and DAWN-FM for the tomography task in Table 2 at noise level $p=5\%$ in the sinogram data for OrganAMNIST and OrganCMNIST datasets [Yang et al., 2023]. For this task, DAWN-FM outperformed DAW-FM on all metrics for both datasets. Some example images for the tomography task for different levels of noise are presented in Figures 10 and 11 in D. In medical image analysis, accurately estimating the size and boundaries of a lobe (such as the one shown in red box in Figure 6) is crucial for diagnostics, especially in cases where its size can influence medical decisions. We quantify uncertainty for both DAW-FM and DAWN-FM by computing mean and standard deviation across 32 samples from the learned posterior. While the mean represents the most probable reconstruction, the standard deviation map highlights regions of higher variability, indicating areas of uncertainty. From Figure 6, DAWN-FM gives a robust understanding of uncertainty by highlighting boundaries of objects as regions of highest uncertainty, a feature similar to what was observed on the MNIST dataset (see Figure 4).
@@ 4 Conclusions
In this paper, we presented DAWN-FM, a framework for addressing highly ill-posed inverse problems by efficiently incorporating data and noise into the FM process. Our experiments showed that our proposed method consistently outperformed existing methods in tasks like image deblurring, with significant improvements in key performance metrics. The ability to sample from the learned posterior enables the exploration of the solution space and facilitates in uncertainty quantification, which is critical for real-world applications. Future work will focus on refining the model to enhance efficiency and applicability to more ill-posed problems, including integration of advanced noise modeling techniques for extreme noise conditions.
@@ Appendix C Experimental Settings
For the image deblurring task, the experiments were conducted on MNIST, STL10 and CIFAR10 datasets. The networks used were Diffusion model, InverseUNetODE as baselines and our proposed methods DAW-FM and DAWN-FM. The key details of the experimental setup for each of these methods are summarized in Tables 3, 4 and 5. For the tomography task, the experiments were conducted using DAW-FM and DAWN-FM methods on OrganAMNIST and OrganCMNIST datasets derived from the MedMNIST data library [Yang et al., 2023]. The key details of the experimental setup are summarized in Table 6. All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch [Paszke et al., 2017].
Table 3: Experimental details for the image deblurring task for our methods DAW-FM and DAWN-FM.
@@
DAW-FM: UNet with 3 levels with [$c$, 16, 32] filters with residual
@@
DAWN-FM: UNet with 3 level levels with [$c$, 16, 32] filters with
@@
DAW-FM: UNet with 5 levels with [3, 16, 32. 64, 128] filters
@@
DAWN-FM: UNet with 5 levels with [3, 16, 32, 64, 128] filters
@@
DAW-FM: max_epochs = 3000;
DAWN-FM: max_epochs = 3000
@@
Table 6: Experimental details for the tomography task for our methods DAW-FM and DAWN-FM.
@@
DAW-FM: UNet with 5 levels with [3, 16, 32. 64, 128] filters
@@
DAWN-FM: UNet with 5 levels with [3, 16, 32, 64, 128] filters
@@
DAW-FM: max_epochs = 3000;
DAWN-FM: max_epochs = 3000